OpenStudy (anonymous):

What is the maximum speed at which a car can travel along a level road when its engine is developing 24kW and there is a resistance to motion of 800N?

OpenStudy (anonymous):

@sauravshakya @Hero @experimentX @Vincent-Lyon.Fr

OpenStudy (anonymous):

by level road you mean no angle ?

OpenStudy (anonymous):

yes horizontal ground

OpenStudy (anonymous):

since power is given by F (force) times V and the force applied is 800 (to equate the resistance) we have 800 * v = 24 * 1000 v = 24 * 1000 / 800 v = 30 m/s

OpenStudy (anonymous):

but isnt there any acceleration?

OpenStudy (anonymous):

maximum speed hence no more acceleration

OpenStudy (anonymous):

thanks alot :)

OpenStudy (anonymous):

yw

OpenStudy (vincent-lyon.fr):

Coolsector has already given the right answer.

OpenStudy (anonymous):

can u help with the question i posted right now wait il tag u